# Integrating representative and non-representative survey data for efficient inference

### v2

@@ 4.4 Inclusion of biased surveys quantified in terms of increased unbiased sample size
One way to measure the reduction in uncertainty, is to measure the number increase of iid samples in the unbiased sample that would be required to obtain the same level of uncertainty. To measure this, we fit the method with the only the Axios-Ipsos data, and then compare the width of the credible intervals of the Axios-Ipsos survey against that of combining the Axios-Ipsos and one of the other surveys, and then ultimately both surveys. The comparison is only made for time-points where an Axios-Ipsos survey is taken. We proceed by calculating $n$ based on the classical frequentist confidence intervals.
@@ 5.2 Selecting a model for $\phi$
In this paper we have provided three possible models for how $\phi$ may evolve over time: the constant model, the linear model, and the random walk model. The choice of selecting the model for $\phi$ depends on the prior belief about the evolution of the bias and the bias variance trade-off of mis-specifying the model. In most applications the bias would neither be constant nor linear, but interpolating the crude error of a random walk with a linear model for $\phi$ when there are few time-points is a reasonable choice. However, as shown in Figure 2, choosing a random walk model is a conservative and surprisingly efficient choice. There is also the possibility of extending the complexity of the model for $\phi$ if the number of time-points is large enough. We could consider employing a moving average model or adding other higher order auto-regressive terms. Yet, we must also be careful about adding too much complexity to a latent parameter, but the possibility exists. Furthermore, we have only considered fitting the same model for both of the non-representative surveys. There is an argument to be made to either specify different models for $\phi_{k}$ for each $k$, or to force the non-representative surveys to share the same jump or slope. In the latter case, this assumes the bias $\phi$ is not necessarily a parameter of the survey, but a parameter of the population that surveys sample from. In either case, the most conservative model (random walk) shown in this paper can be relied upon, having been shown to have a lower MSE than using the unbiased survey, even when mis-specified.
@@ 5.3 Problems with identifying $\phi$
We briefly mentioned in the simulation section that one can run into an issue if $Y_{kt}=n_{kt}$ for at least one $k$ or $t$ in the biased surveys. Suppose that $Y_{kt}=n_{kt}$ for at least one $k$, $t$ in the biased surveys, and $Y_{kt}\leq n_{kt}$ for the unbiased survey(s). Consequently, $\phi_{kt}$ may lie anywhere in $[1,\infty)$, and so the posterior estimate of $\phi_{kt}$ would revert to something resembling the prior. The non-central hyper-geometric model (3) cannot be used if $\phi_{kt}$ is unidentifiable for any $k,t$, as JAGS will not run. It is unclear as to why JAGS cannot handle these cases, as the binomial approximation (4) retains the odds ratio interpretation and runs as expected with reasonable results. This discrepancy is likely related to the complex probability mass function of the non-central hyper-geometric distribution. The question remains: does the unidentifiable $\phi$ produce excessive problems in inference or now-casting? By exclusively analyzing the simulation results where $\phi$ is unidentifiable, there is no obvious change in the bias, and the large reduction in variance is still present. Yet this may be due to the correct specification of the priors. In practice, we would hope that there is a rarely a case where surveys have such a large selection bias where the values of $Y_{kt}$ are so extreme.
@@ 5.4 Extensions to more general data
The proposed method in this paper is only suited for a binary response. However, extensions to a continuous response would be straightforward. The key idea is that we track how the survey’s response bias changes over time. We need not restrict ourselves to any specific type response type, but instead rely on the central limit theorem. For instance, suppose we are interested in surveying salaries and it is presumed that individuals may over-represent their salaries. In this case, a given survey’s estimate may be log normal. Let a given survey $k$ estimate at time $t$ be $\bar{x}_{kt}$, then the sampling distribution of survey estimates could be $\bar{X}_{kt}|\mu_{t},\sigma^{2},\phi_{kt}\sim N(\mu_{t}+\phi_{kt},\sigma^{2}(1-\frac{n_{kt}}{N}))$. Here $\phi_{kt}$ no longer represents an odds ratio, but rather an additive bias that a survey receives from selection bias and/or biased responses.
5.5 The take away
Ultimately, we provide a framework for combining non-representative and representative surveys to reduce uncertainty in survey estimates provided an unbiased survey is present. We can achieve the desirable properties of an unbiased sample, while simultaneously taking advantage of the numerous online surveys that have a large sample size, but suffer from selection bias. These conceptual properties are shown to hold if the movement of selection bias over time is complex. We’ve also provided a model for $\phi_{kt}$ that reliably reduces uncertainty in estimates, even under mis-specification. The framework can be extended to incorporate multiple types of data, and it makes no assumptions on weighting procedures which leaves the door open to practitioners who prefer to weight before making inferences.

### v3

@@ 4.4 Inclusion of biased surveys quantified in terms of increased unbiased sample size
One way to measure the reduction in uncertainty is to measure increase in iid sample size in the unbiased sample required to obtain the same level of uncertainty as the synthesis method. To measure this, we fit the method with the only the Axios-Ipsos data, and then compare the width of the credible intervals of the Axios-Ipsos survey against that of combining the Axios-Ipsos and one of the other surveys, and then ultimately both surveys. The comparison is only made for time-points where an Axios-Ipsos survey is taken. We proceed by calculating $n$ based on the classical frequentist confidence intervals.
@@ 5.2 Selecting a model for $\phi$
In this paper we have provided three possible models for how $\phi$ may evolve over time: the constant model, the linear model, and the random walk model. The choice of selecting the model for $\phi$ depends on the prior belief about the evolution of the bias and the bias variance trade-off of over-parametrizing or under-parameterizing the model. In most applications the bias would neither be constant nor linear, but interpolating the crude error of a random walk with a linear model for $\phi$ when there are few time-points is a reasonable choice. However, as shown in Figure 2, choosing a random walk model is a conservative and surprisingly efficient choice. There is also the possibility of extending the complexity of the model for $\phi$ if the number of time-points is large enough. We could consider employing a moving average model or adding other higher order auto-regressive terms. Yet, we must also be careful about adding too much complexity to a latent parameter, but the possibility exists. Furthermore, we have only considered fitting the same model for both of the non-representative surveys. There is an argument to be made to either specify different models for $\phi_{k}$ for each $k$, or to force the non-representative surveys to share the same jump or slope. In the latter case, this assumes the bias $\phi$ is not necessarily a parameter of the survey, but a parameter of the population that surveys sample from. In either case, the most conservative model (random walk) shown in this paper can be relied upon, having been shown to have a lower MSE than using the unbiased survey, even when over-parametrized.
@@ 5.3 Problems with identifying $\phi$
We briefly mentioned in the simulation section that one can run into an issue if $Y_{kt}=n_{kt}$ for at least one $k$ or $t$ in the biased surveys. Suppose that $Y_{kt}=n_{kt}$ for at least one $k$, $t$ in the biased surveys, and $Y_{kt}\leq n_{kt}$ for the unbiased survey(s). Consequently, $\phi_{kt}$ may lie anywhere in $[1,\infty)$, and so the posterior estimate of $\phi_{kt}$ would revert to something resembling the prior. The non-central hyper-geometric model (3) cannot be used if $\phi_{kt}$ is unidentifiable for any $k,t$, as JAGS will not run. It is not completely understood why JAGS cannot handle these cases, as the binomial approximation (4) retains the odds ratio interpretation and runs as expected with reasonable results. This discrepancy is likely related to the complex probability mass function of the non-central hyper-geometric distribution. The question remains: does the unidentifiable $\phi$ produce excessive problems in inference or now-casting? By exclusively analyzing the simulation results where $\phi$ is unidentifiable, there is no obvious change in the bias, and the large reduction in variance is still present. Yet this may be due to the correct specification of the priors. In practice, we would hope that there is a rarely a case where surveys have such a large selection bias where the values of $Y_{kt}$ are so extreme.
@@ 5.4 Extensions to more general data
The proposed method in this paper is only suited for a binary response. However, extensions to a continuous response would be straightforward. The key idea is that we track how the survey’s response bias changes over time. We need not restrict ourselves to any specific type response type, but instead rely on the central limit theorem. For instance, suppose we are interested in surveying salaries and it is presumed that individuals may over-represent their salaries. In this case, a given survey’s estimate may be log normal. Let a given survey $k$ estimate at time $t$ be $\bar{x}_{kt}$, then the sampling distribution of survey estimates could be $(\bar{X}_{kt}|\mu_{t},\sigma^{2},\phi_{kt})\sim N(\mu_{t}+\phi_{kt},\sigma^{2}(1-\frac{n_{kt}}{N}))$. Here $\phi_{kt}$ no longer represents an odds ratio, but rather an additive bias that a survey receives from selection bias and/or biased responses.
5.5 The take-away
Ultimately, we provide a framework for combining non-representative and representative surveys to reduce uncertainty in survey estimates provided an unbiased survey is present. We can achieve the desirable properties of an unbiased sample, while simultaneously taking advantage of the numerous online surveys that have a large sample size, but suffer from selection bias. These conceptual properties are shown to hold if the movement of selection bias over time is complex. We’ve also provided a model for $\phi_{kt}$ that reliably reduces uncertainty in estimates, even when over-paramterizing. The framework can be extended to incorporate multiple types of data, and it makes no assumptions on weighting procedures which leaves the door open to practitioners who prefer to weight before making inferences.